P3.22Statement of Cash Flow. Presented below are the financial statements for the Amphlett Corporation, as of year-end 2016 and 2015 AMPHLETT CORPORATION Consolldated Balance Sheets As of Year-End (S thousands) 2016 2015 Assets Current 200,000 440,000 615,000 Inventory 600,000 Noncurrent Long-term i 310,000 390,000 (500,000) (500,000 600,000 105,000 95,000 Llablltles & Shareholders Equity 190,000 1.090.000 850,000

As of Year-End ($ thousands) 2016 2015 Shareholders equity 380,000 865,000 300,000 415,000 Retained earni ngs . . . AMPHLETT CORPORATION Consolldated Income Statement For Year Ended ( thousands) 2016 2015 Gross margin... Operating income Net income before taxes . _. . . . ._ . .. . . .. . . 500,000 930,000 150,000 780,000 70,000 850,000 50,000 420,000 780,000 120,000 660,000 . . . 660,000 11,000 . . . . .. .. . . ..

1. Property and equipment costing S450 million was sold for its book value of $200 million 2. Long-term investments were sold for S150 million, which included a gain of S70 million Required Using the above financial data for the Amphlett Corporation, prepare the firms statement of cash flow for 2016 using the indirect method. What does the companys statement of cash flow reveal about the businesss financial health?
Answer #1

Answer:

(all amounts are in thousand $)

Accumulated depreciation of property and equipment sold = Gross value of property and equipment sold - Book value

= $450,000 - $200,000 = 250,000

Depreciation for 2016 = Accumulated depreciation 2016 - (Accumulated depreciation 2016 - Accumulated depreciation of Property and equipment sold)

= 500000 - (500000 - 250,000)

=$250,000

Amortization for 2016 = Intangibles net in 2015 - Intangibles net in 2016 = 105000 - 95000 = 10000

Hence depreciation and amortization for 2016 = 250000 + 10000 = $265,000

Purchase of property and equipment in 2016 =

Gross property and equipment in 2016 - (Gross property and equipment in 2015 - Gross value of property and equipment sold in 2016)

= 1,800,000 - (1,100,000 - 450,000)

=1,150,000

Proceeds from issue of new common stock = Common stock in 2016 + Additional paid-in capital in 2016 - (Common stock in 2015 + Additional paid-in capital in 2015)

= 775,000 + 380,000 - (675,000 + 300,000) = 180,000

Dividend paid in 2016 = Retained earnings in 2015 + Net income in 2016 - Retained earnings in 2016

= 415000 +800000 - 865000

= $350,000

AMPHLETT CORPORATION Statement of Cash Flows For the year ended December 31, 2016 Operating Activities Net Income Adjustment to convert net income to a cash basis Add, Depreciation and amortization Less, Gain on sale of investment Add, Decrease/(Increase) in Account Receivables Add, Decrease/(Increase) in Invento Add, Increase/(Decrease) in Account payable Amount ($thousands $800,000 $260,000 $70,000) ($150,000) $15,000 $50,000 Cash flow from Operating Activities $905,000 Investing Activities Purchase of Pro Proceeds from sale of property and eguipment Purchase of Marketable securities Proceeds from Long Term Investment Cash flow from Investing Activities ($1,150,000) $200,000 ($200,000) $150,000 and Equipment 1,000,000) Financing Activities Proceeds from short term bank debt Proceeds from new issue of common stock Dividend paid Cash flow from Financing Activities $190,000 $180,000 ($350,000) $20,000 Change in Cash flow S75.000) Beginning Cash Balance Ending Cash Balance $90,000 $15,000

Comments on cash flow:

1. Although company has positive cash flow operations but the same is used in purchase of property and equipment. As sales are growing, purchase of property may be to support operations

2. The net cash flow in 2016 is negative.

3. The company has taken short term bank loan and purchased marketable securities.

4. There is increase in account receivables which is higher than proportionate increase in revenue.